You sold and apple for 11.25 each, and you sold a total of 53 apples. However, the profit are not all yours. You have 28% for your pocket and the rest for your sister. So how much will you earn? And your sisters?


1
Expert's answer
2021-08-29T17:28:15-0400

Total amount sold =11.25×53 apples

Amount =596.25

You will earn

"\frac {28}{100} ×596.25 =166.95"

Your sister got the following amount: "596.25-166.95=429.30"